OpenStudy (anonymous):

What are two similarities and two differences between the electromagnetic force and the weak nuclear force?

OpenStudy (anonymous):

Hey I gave you the forces to pick from, you did I see, but don't expect people to do all your work for you, just do a little research, here read this: http://hyperphysics.phy-astr.gsu.edu/hbase/forces/funfor.html

OpenStudy (michele_laino):

I think that the some differences are: 1) the mediator of the electromagnetic force is the photon which is a massless particle, whereas the mediators of the weak force are three bosons, namely 2 charged bosons W, and one neutral boson Z. Now those bosons have a very large mass: \[\begin{gathered} {M_W} = 82Gev/{c^2} \hfill \\ {M_Z} = 92Gev/{c^2} \hfill \\ \end{gathered} \] 2) the coupling constant of electromagnetic force is: \[{g_e} = \sqrt {4\pi \alpha } ,\quad \alpha = \frac{{{e^2}}}{{\hbar c}}\] whereas the coupling constant of theweak force is: \[{g_w} = \sqrt {4\pi {\alpha _w}} \] where \[{{\alpha _w}}\] is the analogous of \[\alpha \] of the electromagnetic theory, furthermore, for the decay of the muon, we have: \[{g_w} = 0.66,\quad {\alpha _w} = \frac{1}{{29}}\] the similarities of those forces, can be thes ones: 1) both forces can be described mathematically using the Feynmann rules, using the famous DIrac matrices and Feynmann diagrams 2) those forces, namely the electromagnetic and the weak force, has been unified by the electroweak theory, developed by these physicists: S. Glashow, A. Salam, S. Weinber Furthermore, an important contributed to this theory was given by the physicist N. Cabibbo, namely his famous Caqbibbo's angles, which are related to the neutral vector and axial vector coupling constants

OpenStudy (michele_laino):

Cabibbo*